# Out-of-Distribution Runtime Adaptation with Conformalized Neural Network Ensembles

### v1

@@ Abstract
We present a method to integrate real-time out-of-distribution (OOD) detection for neural network trajectory predictors, and to adapt the control strategy of a robot (e.g., a self-driving car or drone) to preserve safety while operating in OOD regimes. Specifically, we use a neural network ensemble to predict the trajectory for a dynamic obstacle (such as a pedestrian), and use the maximum singular value of the empirical covariance among the ensemble as a signal for OOD detection. We calibrate this signal with a small fraction of held-out training data using the methodology of conformal prediction, to derive an OOD detector with probabilistic guarantees on the false-positive rate of the detector, given a user-specified confidence level. During in-distribution operation, we use an MPC controller to avoid collisions with the obstacle based on the trajectory predicted by the neural network ensemble. When OOD conditions are detected, we switch to a reachability-based controller to guarantee safety under the worst-case actions of the obstacle. We verify our method in extensive autonomous driving simulations in a pedestrian crossing scenario, showing that our OOD detector obtains the desired accuracy rate within a theoretically-predicted range. We also demonstrate the effectiveness of our method with real pedestrian data. We show improved safety and less conservatism in comparison with two state-of-the-art methods that also use conformal prediction, but without OOD adaptation.
I Introduction
@@
Robotic autonomy leverages learning-based models for perception, trajectory prediction, and control. However, in many situations, the performance of these models depends highly on the distribution of the data used in inference. Specifically, learned models generally exhibit satisfactory performance when the input data for inference comes from the distribution of the training data (i.e., the in-distribution setting), with a degradation in performance when the distribution of the input data for inference differs from the training data distribution (i.e., the out-of-distribution (OOD) setting). This degradation in performance in OOD settings has limited the utilization of deep-learning models in safety-critical applications. In this work, we introduce a safe adaptive controller for autonomous robots, which enables robots to operate safely in their environments, across scenarios with in-distribution and out-of-distribution behavior by dynamic agents operating alongside these autonomous robots.
Figure 1: Our algorithm enables an autonomous vehicle (the red vehicle driving from left to right) to navigate safely, with provable guarantees, while avoiding collisions with a pedestrian (cyan marker halfway across the crosswalk, moving from top to bottom), even in out-of-distribution settings, by composing a model of pedestrian behavior with reachability sets for the pedestrian in an MPC-based controller. (The larger black circle represents the reachable set of the pedestrian.)
A neural network ensemble—a collection of neural networks trained on the same data—provides a simple and effective means of measuring the uncertainty of a neural network prediction without knowledge of the underlying probability distribution of the training data. Generally, these networks tend to agree on their predictions for inputs similar to the training data, and disagree more widely as the input departs farther from the training distribution, providing a means for uncertainty quantification. Unfortunately, measures of disagreement for the ensemble are not statistically calibrated signals—they are not tied to an explicit probability of the data being in or out-of distribution. We therefore use the framework of conformal prediction to calibrate the ensemble disagreement, so that it can be used to rigorously quantify the probability that a new input is out-of-distribution.
Our proposed control scheme uses model predictive control (MPC) accounting for collision avoidance, with obstacle prediction when operating in-distribution and guaranteed-safe reachability sets when operating out-of-distribution. In our work, we utilize conformal prediction to calibrate out-of-distribution detection through a neural network ensemble, which ultimately determines the control strategy applied within our controller.
We demonstrate our control method in an autonomous driving scenario where an autonomous vehicle (AV) must avoid colliding with a pedestrian crossing the road, illustrated in Figure 1. We utilize a trajectory prediction model trained from real pedestrian data [47]. However, to demonstrate the safety guarantees provided by our controller in the OOD setting, we modify the real data such that the pedestrian occasionally behaves differently from the training data, by running head-on towards the approaching car, simulating a worst-case scenario. The objective of the car is to avoid colliding with the pedestrian (in both nominal and OOD regimes), while still reaching a goal position on the road, if possible. In our evaluations, we find that our method reliably detects out-of distribution behavior by the pedestrian at the rate predicted by the conformal prediction analysis, and goes past the pedestrian, when possible, without any collisions in all cases, unlike other existing methods. Although the SPDE control method [22] avoids collisions with the pedestrian in the nominal scenario, SPDE results in collisions in the OOD regime. Further, we note that the MARC method [27] tends to be overly conservative, failing to go past the pedestrian to reach the goal although it avoids collisions with the pedestrian in all cases.
@@
  • We derive a model predictive control architecture for learned predictive models that safely adapts to out-of-distribution model predictions. Our controller combines an efficient MPC-based controller for in-distirbution settings with a more conservative reachable-set-based controller in out-of-distribution settings.
  • In addition, we introduce an out-of-distribution detector with a guaranteed error rate, based on conformal prediction, utilized within the switching procedure in our control framework to provide safety assurances in both in-distribution and out-of-distribution domains.
II Related Works
II-A Out-of-Distribution Detection in Neural Networks
Neural networks tend to produce accurate predictions when the input data is within the distribution of the training data, and in many situations, are overconfident when generating highly inaccurate predictions with out-of-distribution inputs [15, 28], spurring the development of strategies for out-of-distribution detection. Many existing methods define measures of uncertainty used in detecting OOD inputs based on the outputs of the pertinent neural network [23, 46], such as measures based on the magnitude of the softmax probabilities in the output layer in classification problems [15] and uncertainty measures based on the weights and activation patterns of neurons in the hidden layers [38, 10, 37, 33]. These methods offer the benefit of not requiring additional training of the associated neural networks, enabling post-training utilization. Other methods incorporate additional branches to the underlying network architecture, which estimate the network’s confidence in its predictions [7, 20, 14, 39]. These methods further modify the loss function used in training, to encourage the network to estimate distinguishable confidence scores between in-distribution and out-of-distribution settings. Other methods leverage network ensembles to estimate the confidence of the ensemble in its predictions and accordingly detect out-of-distribution inputs [26]. We refer interested readers to [48] for a detailed survey on out-of-detection methods in deep learning. Although these existing methods work well in many situations, these methods do not provide provable OOD detection probabilistic guarantees, in general.
II-B Conformal Prediction
The theory of conformal prediction provides a statistical method for generating prediction regions from online observations and/or predictions, with minimal assumptions on the probability distribution from which the observed data is sampled [45, 32]. The resulting prediction regions are associated with a specified significance level, which are valid probabilistically: for a specified significance level $\alpha$, the associated prediction region contains the target prediction ${(1-\alpha)\%}$ of the time. Since conformal prediction is agnostic to the prediction method, conformal prediction has been applied to a broad variety of machine learning approaches, with applications in automated decision support systems [35, 3]; medical diagnostics, prognostics, and drug discovery [29, 11, 1]; and safe trajectory prediction and motion planning [25, 22, 9, 21, 36], which we discuss in greater detail in the next subsection. In contrast to other uncertainty quantification methods, conformal prediction enables uncertainty quantification without making assumptions on the shape of the data-generating distribution.
II-C Safe Trajectory Prediction and Motion Planning
Motion planning in the presence of dynamic agents has been extensively studied. To simplify the added complexity from the dynamic nature of the environment and improve efficiency, some existing methods consider dynamic agents as being static at each time step, constantly replanning at each time step [41, 16]. These methods trade completeness and optimality for efficiency. In contrast, other methods predict the trajectory of dynamic agents, explicitly considering the predicted trajectories of these agents in computing a safe motion plan via graph-search methods or sampling-based methods [34, 30, 43]. However, these methods make assumptions on the behavior of the dynamic agents [42], which might not hold in practice, and do not provide probabilistic guarantees on the safety of the resulting motion plans. To provide safety guarantees in trajectory prediction, some existing methods leverage conformal prediction to compute prediction regions for an associated nonconformity measure, with applications to the detection of unsafe situations in early-warning systems [25], verification of autonomous systems [12, 8], and motion planning [5, 22, 27]. In our work, we utilize conformal prediction to provide probabilistic guarantees on the safety of a trajectory prediction module and model predictive controller for autonomous navigation of an agent. In contrast to [22] and [27], which apply conformal prediction within a model predictive control framework and reachable-set-based control scheme, respectively, we introduce an adaptive control scheme that provides probabilistic safety assurances, without being overly conservative.
III Problem Formulation
In this paper, we consider an autonomous robot operating in an environment with other agents, where the dynamics model of the robot is given by:
$\mathbf{x}_{t+1}=f_{e}(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1)
@@ III Problem Formulation
where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{e}}$ represents the state of the robot at time step $t$, $\mathcal{X}$ represents its workspace, and $\mathbf{x}_{0}\in\mathcal{X}$ represents its initial state. The control input at time step $t$ is represented by $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$, where the set of permissible control inputs is represented by $\mathcal{U}$. The function ${f_{e}:\mathbb{R}^{n_{e}}\times\mathbb{R}^{m}\to\mathbb{R}^{n_{e}}}$ describes the robot’s dynamics, which are presumed known.
@@
The robot seeks to navigate to a desired goal location optimally without colliding with other agents in its environment. We do not assume that these dynamic agents cooperate with the robot, allowing for potential adversarial behaviors by these agents. Let the random variable $Y$ denote the trajectory of an agent. We assume that $Y$ follows an unknown stationary distribution $\mathcal{D}$ over the trajectories of agents in the environment, i.e., ${Y\sim\mathcal{D}}$, where ${Y_{t}\in\mathbb{R}^{n_{a}}}$ denotes the state of the environment at time $t$.
We use $y$ when referring to a realization of one of these trajectories, with $y_{t}$ referring to the observed environment state at time $t$. The dynamics of these agents are a priori unknown, but can be approximated with a learned model $f_{a}:\mathbb{R}^{n_{a}}\times\mathbb{R}^{m}\to\mathbb{R}^{n_{a}}$, based on a dataset $D$ where $D:=\{y^{(1)},\dots,y^{(k)}\}$, and $y^{(i)}$ represents the $i$th complete observed trajectory of the agents in the environment.
Assumption 1.
@@
Problem 1 (Safe Adaptive Controller).
Given the robot’s dynamics in (1), and a set of observed in-distribution trajectories of agents $D:=\{y^{(1)},\dots,y^{(k)}\}$, design a safe adaptive controller that enables an autonomous robot to safely navigate its environment, when operating alongside other agents whose behavior (trajectories) may be “in-distribution” or “out-of-distribution.”
IV Model Predictive Control
To navigate safely in its environment, an autonomous robot must be able to predict the future trajectory of other agents in its environment, for collision avoidance.
Consequently, given the dataset $D$ of trajectories of the dynamic agents in the environment of the autonomous robot, we seek to learn a predictive model $f_{a}$ that predicts the state of the dynamic agents in the environment at the next time step when queried at the current time step $t$. We note that many approaches exist for trajectory prediction: notably, deep-learned models such as feed-forward neural networks, recurrent neural networks, and transformers could be utilized for trajectory prediction. Further, the trajectory prediction task can be accomplished using a single predictive model or using a group of models, referred to as an ensemble.
In this work, we utilize a neural network ensemble, consisting of $n$ individual multi-layer perceptron (MLP) neural networks. Neural network ensembles enable estimation of the probability distribution of the outputs of the model, which is valuable in preventing overfitting and in confidence (covariance) estimation [13]. We note that approaches such as Bayesian neural networks [17] and dropout methods for neural networks [18] can be considered as special examples of ensemble learning, providing similar advantages. We utilize relatively small MLPs as the individual models in the ensemble to demonstrate the effectiveness of our algorithm with relatively simple predictors. Figure 2 depicts the architecture of each neural network in the neural network ensemble for trajectory prediction. In addition, we note that the individual models comprising the ensemble are not required to have the same architecture.
@@ IV Model Predictive Control
${\mathbf{s}}_{t+1}=\frac{1}{n}\sum_{i=1}^{n}f_{a}^{i}(\mathbf{s}_{t-\ell:t}),$ (2)
$\Sigma_{t+1}=\frac{1}{n-1}\sum_{i=1}^{n}(f_{a}^{i}(\mathbf{s}_{t-\ell:t})-{\mathbf{s}}_{t+1})(f_{a}^{i}(\mathbf{s}_{t-\ell:t})-{\mathbf{s}}_{t+1})^{\intercal},$ (3)
where $f_{a}^{i}$ represents the $i$th model in the ensemble, ${\mathbf{s}}_{t+1}$ denotes the mean prediction of the state of the environment at time step $t+1$ (with the prediction made at time $t$), and ${\Sigma}_{t+1}$ represents the unbiased estimate of the corresponding covariance matrix, for the predictions made by the models in the ensemble at time step $t$ (with the input to the network given by $\mathbf{s}_{t-\ell:t}$, a concatenation of the states of the environment over the preceding ${\ell}$ steps including the current time step).
We begin with designing a controller for the nominal case, where the trajectories of the dynamic agents follow the distribution given by $\mathcal{D}$. In this case, we design a model predictive control-based controller leveraging the predictive model $f_{a}$ to enforce collision-avoidance constraints. The corresponding model predictive control (MPC) problem is given by:
@@
$\mathrm{s.t.}$ $\mathbf{x}_{\tau+1}=f_{e}(\mathbf{x}_{\tau},\mathbf{u}_{\tau})\enspace\forall\tau=1,\dots,T-1$
$\mathbf{s}_{\tau+1}=f_{a}(\mathbf{s}_{\tau})\enspace\forall\tau=1,\dots,T-1$
$g(\mathbf{x}_{\tau},\mathbf{s}_{\tau})\leq 0\enspace\forall\tau=1,\dots,T,$
where $J:\mathbb{R}^{T\cdot n_{e}}\times\mathbb{R}^{T\cdot m}\to\mathbb{R}$ denotes the objective function, and $\mathbf{x}_{1:T}$ and $\mathbf{u}_{1:T}$ denote the concatenation of the optimization variables representing the robot’s states and control inputs over the time steps $1,\dots,T$, respectively. We assume that the MPC problem has a planning horizon of $T$ time steps. We specify constraints enforcing initial conditions, collision avoidance, and other undesirable interactions in $g:\mathbb{R}^{n_{e}}\times\mathbb{R}^{n_{a}}\to\mathbb{R}$. The MPC problem is periodically re-solved by the controller over the course of the simulation, with a new solution computed every $H$ time steps: after solving an instance of the optimization problem, the ego vehicle applies the first $H$ control inputs $u_{t:t+H}$ over the next $H$ time steps, and subsequently solves a new instance of the MPC problem with updated observations.
However, we note that the resulting MPC-based controller (4) fails to provide safety guarantees, particularly when the dynamic agents exhibit out-of-distribution behavior. To address this limitation, we develop an adaptive controller, considering both in-distribution and out-of-distribution settings. Our out-of-distribution controller leverages the reachable set of the dynamic agents, representing the set of states that the dynamic agents can reach over a specified time duration. We denote the reachable set for the agents in the environment at time step $t$ by $\mathcal{R}_{t}\subseteq\mathcal{S}$, which we compute from:
$\mathcal{R}_{t+1}=\textsc{Reach}(\mathcal{R}_{t}),$ (5)
with $\mathcal{R}_{1}=\{\mathbf{s}_{1}\}$, where ${\textsc{Reach}:\mathbb{S}\to\mathbb{S}}$ and $\mathbb{S}$ denotes all subsets of $\mathbb{R}^{n_{a}}$. In computing the reachable set of the dynamic agents, we make the following assumption:
Assumption 2.
We assume that a maximum speed ${v_{max}\in\mathbb{R}}$ exists, which all agents in the environment of the autonomous robot do not exceed.
This assumption is realistic, as the velocity of dynamic agents such as pedestrians can be bounded from above. To provide safety assurances in out-of-distribution settings, we consider an MPC problem where the predictive model for the state of the agents in (4) is replaced with a reachable-set-based constraint. The corresponding MPC problem is given by:
@@
$\mathrm{s.t.}$ $\mathbf{x}_{\tau+1}=f_{e}(\mathbf{x}_{\tau},\mathbf{u}_{\tau})\enspace\forall\tau=1,\dots,T-1$
$\mathcal{R}_{\tau+1}=\textsc{Reach}(\mathcal{R}_{\tau})\enspace\forall\tau=1,\dots,T-1$
$g(\mathbf{x}_{\tau},\mathcal{R}_{\tau})\leq 0\enspace\forall\tau=1,\dots,T,$
@@
To design a safe, OOD-adaptive controller for Problem 1, we compose the MPC-based controller for the nominal setting (MPC I) with the reachable-set-based controller (MPC II) to enable safe navigation by autonomous robots, necessitating the design of a rule for switching between both controllers. Specifically, an autonomous robot must be able to distinguish between dynamic agents which are acting nominally from those agents with out-of-distribution behavior, concepts which we precisely define later in this work. To address this challenge, we introduce an out-of-distribution detector, which leverages conformal prediction to provide probabilistic guarantees. We state the out-of-distribution detection problem, prior to discussing our proposed approach to solving this problem:
Problem 2 (OOD detection).
@@
V Out-of-Distribution Detection
We detect out-of-distribution samples based on measures of uncertainty derived from the empirical covariance matrix ${\Sigma}_{t+1}$, defined in (3). Many measures of uncertainty exist in ensemble learning, such as the spectral norm (also known as the $2$-norm), the Frobenius norm, and the trace of the empirical covariance matrix. In this work, we use the spectral norm as our uncertainty measure:
$\rho_{t+1}=\|{\Sigma}_{t+1}\|_{2},$ (7)
noting that $\Sigma_{t+1}$ is positive semidefinite and symmetric for all $t$. In addition, we note that other measures of uncertainty could be used, such as [33] and [24]. However, these methods require access to the training dataset, unlike our approach.
We leverage conformal prediction to provide provable probabilistic safety guarantees for the safe adaptive controller introduced in this work. We begin with a brief introduction of conformal prediction.
V-A Conformal Prediction
Conformal prediction was introduced in [45] and [32], to compute prediction regions for complex predictive models without making assumptions about the distribution underlying the predictive model [2]. We present a brief introduction based on the assumption that the observations arise from independent and identical distributions, for easier presentation. However, we note that the theory of conformal prediction only requires the assumption of exchangeability. Observations in a set ${\{\rho^{(1)},\ldots,\rho^{(N)}\}}$ are exchangeable if all possible orderings of these observations are equally likely to occur. Consequently, exchangeability represents a weaker assumption compared to the assumption of independent, identically-distributed observations.
Let $R^{(1)},\dots,R^{(N)}$ be $N$ independent and identically distributed random variables. The variable $R^{(i)}$ is generally referred to as the nonconformity measure or score. Conformal prediction enables us to construct a valid prediction region $C$ associated with a given failure probability ${\delta\in(0,1)}$, satisfying:
$P(R^{(0)}\leq C)\geq 1-\delta,$ (8)
[…]

### v3

@@ Abstract
We present SODA-MPC, a Safe, Out-of-Distribution-Adaptive Model Predictive Control algorithm that uses an ensemble of learned models for prediction with a runtime monitor to flag unreliable out-of-distribution (OOD) predictions. When an OOD situation is detected, SODA-MPC triggers a safe fallback control strategy based on reachability, producing a control framework that achieves the high performance of learning-based models while preserving the safety of reachability-based control. We demonstrate the method in the context of an autonomous vehicle, driving among dynamic pedestrians, where SODA-MPC uses a neural network ensemble for pedestrian prediction. We use the maximum singular value of the empirical covariance among the ensemble as the OOD signal for the runtime monitor. We calibrate this signal using conformal prediction to derive an OOD detector with probabilistic guarantees on the false-positive rate, given a user-specified confidence level. During in-distribution operation, the MPC controller avoids collisions with a pedestrian based on the trajectory predicted by the mean of the ensemble. When OOD conditions are detected, the MPC switches to a reachability-based controller to avoid collisions with the reachable set of the pedestrian assuming a maximum pedestrian speed, to guarantee safety under the worst-case actions of the pedestrian. We verify SODA-MPC in extensive autonomous driving simulations in a pedestrian-crossing scenario. Our model ensemble is trained and calibrated with real pedestrian data, showing that our OOD detector obtains the desired accuracy rate within a theoretically-predicted range. We empirically show improved safety and task completion compared with two state-of-the-art MPC methods that also use conformal prediction but without OOD adaptation. Further, we demonstrate the effectiveness of our method with the large-scale multi-agent predictor Trajectron++, using large-scale traffic data from the nuScenes dataset for training and calibration.
keywords:
Conformal prediction, out-of-distribution (OOD) detection, and ensemble learning.
1 Introduction
@@
Robotic autonomy stacks often leverage learning-based models for perception, trajectory prediction, and control. However, in many situations, the performance of these models depends highly on the distribution of the input data encountered at runtime. Learned models can exhibit strong performance when the runtime input data is similarly distributed to the training data (the in-distribution setting), but performance suffers when the runtime inputs are significantly different from the training data (the out-of-distribution, or OOD, setting [^1] We specifically consider the OOD situation known as covariate shift, where the marginal distribution of the input data changes between training and runtime. In our problem, this arises from observed agents expressing unusual behavior not well-represented in the training data.) (Nguyen et al., 2015). This deterioration in performance in OOD settings has limited the utilization of deep-learning models in safety-critical applications. In this work, we introduce a Safe OOD-Adaptive Model Predictive Controller (SODA-MPC), which enables robots to operate safely while sharing their task space with other agents that exhibit both in-distribution and OOD behavior.
Figure 1: The SODA-MPC algorithm avoids pedestrian collisions in both the nominal (left) and adversarial (right) behavior settings. The algorithm correctly detects the adversarial pedestrian behavior as out-of-distribution, and switches its control strategy to the safe reachable-set-based MPC.
We consider a standard autonomous vehicle (AV) control scheme, which uses an MPC with collision constraints to avoid other agents in the scene. The future trajectories of these agents (required for the collision avoidance constraint in the MPC) are predicted with a learned trajectory prediction model. We address a critical flaw in this architecture: such trajectory predictions are only reliable in in-distribution regimes. Our SODA-MPC algorithm alleviates this problem by introducing a statistically calibrated runtime monitor to detect OOD situations online and switch to a safe fallback controller in those cases. Our runtime monitor uses a neural network ensemble for OOD detection. However, we note that widely-used measures of disagreement for ensemble models are not statistically calibrated signals—they are not tied to an explicit probability of the data being in or out of distribution. Hence, we use conformal prediction to calibrate an OOD detector for the ensemble. When the detector flags an OOD scenario, the MPC ignores the neural network prediction and avoids colliding with the forward reachable sets of the agents, which is a guaranteed-safe fallback controller. The reachable sets are computed assuming a maximum speed bound for the agents.
We demonstrate SODA-MPC in a scenario where an AV must avoid colliding with a pedestrian crossing the road, illustrated in Figure 1. We use a trajectory prediction model ensemble trained from real pedestrian data (Yang et al., 2019). To demonstrate the safety guarantees provided by our controller, we use a handful of held-out trajectories of the original data (which we refer to as nominal), and create modified instances of these so that the pedestrian exhibits OOD behavior by running head-on towards the approaching car, a worst-case scenario for the AV. The car must avoid colliding with the pedestrian in both nominal and OOD regimes while attempting to reach a goal position on the road. In our evaluations, we find that our method reaches the goal position in a majority of cases and reliably detects OOD behavior such that it avoids all pedestrian collisions across all tests. We compare against two recent baselines that also use conformal prediction for AV control with learned pedestrian prediction models: SPDE (Lindemann et al., 2023), which calibrates confidence intervals for the learned model, and MARC (Muthali et al., 2023), which calibrates reachable sets for the learned models. Neither method directly addresses OOD behavior. We find that our algorithm is safer than SPDE in OOD regimes and more efficient than MARC which tends to be overly conservative in in-distribution regimes, as demonstrated in Figure 4. In addition, we demonstrate the success of our detector in identifying OOD scenarios when applied to the large-scale trajectory predictor Trajectron++ (Salzmann et al., 2020).
@@
  • We derive an MPC architecture, SODA-MPC, for learned predictive models that safely adapts to out-of-distribution model predictions. Our MPC switches to a guaranteed-safe reachable set based prediction in out-of-distribution regimes.
  • In addition, we introduce an OOD runtime monitor with a guaranteed error rate using neural network ensembles. The monitor is statistically calibrated with conformal prediction, supplying the switching signal for our SODA-MPC.
2 Related Works
2.1 Out-of-Distribution Detection in Neural Networks
Methods for OOD detection can be broadly grouped into training-free and fine-tuning methods. Training-free methods directly leverage the outputs (Liu et al., 2020; Wang et al., 2021; Hendrycks and Gimpel, 2016), e.g., softmax probabilities, and activation patterns (Sun et al., 2021; Dong et al., 2022; Sun and Li, 2022; Sharma et al., 2021) of a trained model to detect OOD inputs. Fine-tuning methods modify the underlying architecture of the model, including the loss function, to estimate the network’s confidence in its predictions to distinguish between in-distribution and out-of-distribution settings, e.g., (DeVries and Taylor, 2018; Lee et al., 2017; Hein et al., 2019; Thulasidasan et al., 2019; Madras et al., 2019). We refer interested readers to (Yang et al., 2021) for a more detailed discussion. In the context of robotics, Sinha et al. (2024) detects visual anomalies with an LLM, and enacts a safe fallback MPC if a danger is detected. Although these existing methods work well in many situations, they do not provide provable OOD detection probabilistic guarantees, in general.
2.2 Conformal Prediction in Trajectory Prediction and Motion Planning
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) provides a black-box, distribution-free method for generating prediction regions from online observations, making it ideal for wide-ranging applications, e.g., in automated decision support systems (Straitouri et al., 2023; Babbar et al., 2022) and in safe trajectory prediction and motion planning (Luo et al., 2022b; Lindemann et al., 2023; Dixit et al., 2023; Lekeufack et al., 2023; Sun et al., 2023). In trajectory prediction, some existing methods leverage conformal prediction for the detection of unsafe situations in early-warning systems (Luo et al., 2022b), verification of autonomous systems (Fan et al., 2020; Dietterich and Hostetler, 2022), and motion planning (Chen et al., 2021; Lindemann et al., 2023; Muthali et al., 2023). In our work, we utilize conformal prediction to provide probabilistic guarantees on the safety of a trajectory prediction module and model predictive controller for autonomous navigation of an agent. Similar to out work, (Lindemann et al., 2023) and (Muthali et al., 2023) apply conformal prediction to learned trajectory prediction models to obtain in-distribution safety guarantees for an MPC. In contrast, we use conformal prediction and a safe fallback controller to obtain safety for both in-distribution and OOD settings.
3 Problem Formulation
@@
We consider an autonomous robot operating in an environment with other agents, where the dynamics model of the robot is given by: ${\mathbf{x}_{t+1}=f_{e}(\mathbf{x}_{t},\mathbf{u}_{t}),}$ where $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{e}}$ represents the state of the robot at time step $t$, $\mathcal{X}$ represents its workspace, $\mathbf{x}_{0}\in\mathcal{X}$ represents its initial state, ${\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}}$ denotes its control input at $t$, where $\mathcal{U}$ represents the set of permissible control inputs, and ${f_{e}:\mathbb{R}^{n_{e}}\times\mathbb{R}^{m}\to\mathbb{R}^{n_{e}}}$ describes the robot’s dynamics, which are presumed known.
@@
The robot seeks to navigate to a goal location optimally without colliding with other agents in its environment. Our method can handle multiple or single agents. We do not assume that these dynamic agents cooperate with the robot, allowing for potential adversarial behaviors. Let the random variable $Y$ denote the joint state of multiple agents, which follows an unknown stationary distribution $\mathcal{D}$ over the states of agents, i.e., ${Y\sim\mathcal{D}}$, where ${Y_{t}\in\mathbb{R}^{n_{a}}}$ denotes the state of the agents at time $t$, and $n_{a}$ is the stacked state dimension of all the agents. We use $y$ when referring to an observation of one of these multi-agent states, with $y_{t}$ referring to the agents’ state at time $t$. The true dynamics of these agents are unknown, but can be approximated with a learned model $f_{a}:\mathbb{R}^{n_{a}}\times\mathbb{R}^{m}\to\mathbb{R}^{n_{a}}$, learned from a dataset $D$ where $D:=\{y^{(1)},\dots,y^{(k)}\}$, and $y^{(i)}$ represents the $i$th complete observed joint state of the agents in the environment.
Assumption 1
@@
Problem 1 (Safe Adaptive Controller)
Given the robot’s dynamics and a set of observed in-distribution trajectories of agents $D:=\{y^{(1)},\dots,y^{(k)}\}$, design an MPC controller that enables an autonomous robot to safely navigate its environment, when operating alongside other agents whose trajectories may be “in-distribution” or “out-of-distribution.”
4 Model Predictive Control
In this work, we utilize a learned ensemble, consisting of $n$ individual multi-layer perceptron (MLP) neural networks for trajectory prediction, given a dataset $D$. Neural network ensembles enable estimation of the probability distribution of the outputs of the model, which is valuable in preventing overfitting and in confidence estimation (Gustafsson et al., 2020). We note that approaches such as Bayesian neural networks (Jospin et al., 2022) and dropout methods for neural networks (Labach et al., 2019) can be considered as special examples of ensemble learning, providing similar advantages. We use relatively small MLPs as the individual models in the ensemble to demonstrate the effectiveness of our algorithm with simple predictors.
@@ 4 Model Predictive Control
$\intercal\displaystyle{\mathbf{s}}_{t+1}=\frac{1}{n}\sum_{i=1}^{n}f_{a}^{i}(\mathbf{s}_{t-\ell:t}),\quad\Sigma_{t+1}=\frac{1}{n-1}\sum_{i=1}^{n}(f_{a}^{i}(\mathbf{s}_{t-\ell:t})-{\mathbf{s}}_{t+1})(f_{a}^{i}(\mathbf{s}_{t-\ell:t})-{\mathbf{s}}_{t+1})^{\intercal},$ (1)
where $f_{a}^{i}$ represents the $i$th model in the ensemble, ${\mathbf{s}}_{t+1}$ denotes the mean prediction of the state of the environment at time step $t+1$ (with the prediction made at time $t$), and ${\Sigma}_{t+1}$ represents the unbiased estimate of the corresponding covariance matrix for the predictions made by the models in the ensemble at time step $t$. The input to each network is $\mathbf{s}_{t-\ell:t}$, a concatenation of the (observed or predicted) states of the agents over the preceding ${\ell}$ steps, including the current time step. We write ${\mathbf{s}}_{t+1}=f_{a}(\mathbf{s}_{t-\ell:t})$ to denote the ensemble mean.
We begin with designing a controller for the nominal case, where the trajectories of the dynamic agents follow the distribution given by $\mathcal{D}$. In this case, we design a model predictive controller leveraging the predictive model $f_{a}$ to enforce collision-avoidance constraints. The corresponding model predictive control (MPC) problem is given by:
@@
$\mathrm{s.t.}$ $\mathbf{x}_{\tau+1}=f_{e}(\mathbf{x}_{\tau},\mathbf{u}_{\tau}),\enspace\mathbf{s}_{\tau+1}=f_{a}(\mathbf{s}_{\tau-l:\tau}),\enspace\forall\tau=0,\dots,T-1,$
$g(\mathbf{x}_{\tau},\mathbf{s}_{\tau})\leq 0,\enspace\forall\tau=1,\dots,T,$
where $J:\mathbb{R}^{T\cdot n_{e}}\times\mathbb{R}^{T\cdot m}\to\mathbb{R}$ denotes the objective function and $\mathbf{x}_{1:T}$ and $\mathbf{u}_{0:T-1}$ denote the concatenation of the optimization variables representing the robot’s states and control inputs over the pertinent time steps respectively. We assume we have access to $s_{0}$, from which we predict the trajectory $\mathbf{s}_{\tau-l:\tau}$${\forall\tau}$, and $\mathbf{x}_{0}$, the initial state of the robot. We assume that the MPC problem has a planning horizon of $T$ time steps. We specify constraints enforcing initial conditions, collision avoidance, and other undesirable interactions in $g:\mathbb{R}^{n_{e}}\times\mathbb{R}^{n_{a}}\to\mathbb{R}$. The MPC problem is re-solved every $H$ time steps with updated observations.
We note that the resulting MPC-based controller (2) fails to provide safety guarantees, particularly when the dynamic agents exhibit OOD behavior. To address this limitation, we develop an adaptive controller, considering both in-distribution and OOD settings. Our OOD controller leverages the reachable set of the dynamic agents, representing the set of states that the dynamic agents can reach over a specified time duration. We denote the reachable set for the agents in the environment at time step $t$ by $\mathcal{R}_{t}\subseteq\mathcal{S}$, which we compute from: ${\mathcal{R}_{t+1}=\textsc{Reach}(\mathcal{R}_{t})}$, with $\mathcal{R}_{0}=\{\mathbf{s}_{0}\}$, where ${\textsc{Reach}:\mathbb{S}\to\mathbb{S}}$ and $\mathbb{S}$ denotes all subsets of $\mathbb{R}^{n_{a}}$. In computing the reachable set of the dynamic agents, e.g., via velocity-based reachability analysis, we make the following assumption:
Assumption 2
We assume knowledge of a maximum speed ${v_{max}\in\mathbb{R}}$ for all agents.
This assumption is realistic, as we can use reasonable bounds for the top speed of agents in the robot’s workspace, such as pedestrians. To provide safety assurances in OOD settings, we consider an MPC problem where the predictive model for the state of the agents in (2) is replaced with a reachable-set-based constraint. The corresponding MPC problem is given by:
@@
$\mathrm{s.t.}$ $\mathbf{x}_{\tau+1}=f_{e}(\mathbf{x}_{\tau},\mathbf{u}_{\tau}),\enspace\mathcal{R}_{\tau+1}=\textsc{Reach}(\mathcal{R}_{\tau}),\enspace\forall\tau=0,\dots,T-1,$
$g(\mathbf{x}_{\tau},\mathcal{R}_{\tau})\leq 0,\enspace\forall\tau=1,\dots,T,$
@@
To design a safe, OOD-adaptive controller for Problem 1, we compose the MPC-based controller for the nominal setting (MPC I) with the reachable-set-based controller (MPC II) to enable safe navigation by autonomous robots, necessitating the design of a rule for switching between both controllers. Specifically, an autonomous robot must be able to distinguish between dynamic agents which are acting nominally from those agents with OOD behavior, concepts which we precisely define later in this work. To address this challenge, we introduce an OOD detector, which leverages conformal prediction to provide probabilistic guarantees. We state the OOD detection problem, prior to discussing our proposed approach to solving this problem:
Problem 2 (OOD detection)
@@
5 Out-of-Distribution Detection
We utilize the spectral norm $\rho_{t+1}$ of the empirical covariance matrix ${\Sigma}_{t+1}$ (1) as our uncertainty measure to detect OOD samples, with ${\rho_{t+1}=\|{\Sigma}_{t+1}\|_{2}}$, noting that $\Sigma_{t+1}$ is positive semidefinite and symmetric for all $t$. However, related measures such as the trace, spectral norm, and Frobenius norm of the covariance matrix and techniques such as (Sharma et al., 2021) and (Luo et al., 2022a) can also be used as the uncertainty measure. We note that the techniques in (Sharma et al., 2021) and (Luo et al., 2022a) require access to the whole training dataset, while we only require a handful of held-out data samples. We leverage conformal prediction to provide provable probabilistic safety guarantees for the safe adaptive controller introduced in this work (see Appendix A for a brief introduction of conformal prediction).